OpenStudy (anonymous):

How do I raise number to the 1/2 power????

OpenStudy (agreene):

\[a^{1/2}=\sqrt a\] does that help?

OpenStudy (anonymous):

I know that it means sq root. but I actually want to know how to raise a number to a 1/2

OpenStudy (agreene):

you mean, you want to know how to find a square root?

OpenStudy (anonymous):

How would I solve 7 ^1/2??

OpenStudy (agreene):

Well, there are three options. #1: don't. It is irrational thereby meaning 7^(1/2) is in its most exact form. #2: use a calculator with a sqrt button #3: try to estimate it... 2*2=4 which is less than 7 3*3=6 which is more than 7. It can't be more than 7 so we know it is 2.something 2.1*2.1=4.41 2.9*2.9=8.41 which is more so it cant be this one 2.6*2.6= 6.76 thats really close but less than 7. 2.7*2.7= 7.29 too much, must be 2.6 Now we are at 2.6..... 2.64*2.64 = 6.9696 really really close :D 2.65*2.65 = 7.0225 too much... must be 2.64.... keep doing that until you're happy.